OpenStudy (anonymous):

**********medals and fans*********Lina is making a spreadsheet for an assignment. The spreadsheet, when completed, will be 220 lines long, and Lina has already completed 78 lines. x + 78 = 220 How many lines does Lina have left to complete?

OpenStudy (anonymous):

x + 78 = 220 x = 220 - 78 x = 142